OpenStudy (anonymous):

prove the identity sin^2x + tan^2xsin^2x = tan^2x

OpenStudy (anonymous):

Double angle formula

OpenStudy (anonymous):

sin^2(x) (1+tan^2(x)) = sin^2(x)*sec^2(x) = tan^2(x)

OpenStudy (anonymous):

nice, that's an easier way to do it ^ you could also do it this way: tan^2x - sin^2x= tan^2xsin^2x sin^2x/cos^2x - sin^2x= tan^2xsin^2x sin^2x ( 1/cos^2x - 1) = tan^2xsin^2x sin^2x ( 1-cos^2x/ cos^2x) = tan^2xsin^2x sin^2 x (sin^2x / cos^2x) = tan^2xsin^2x sin^2xtan^2x = tan^2xsin^2x

OpenStudy (anonymous):

thanks but any without sec

OpenStudy (gw2011):

sec x = 1/cos x Therefore, sec^2x = 1/cos^2x and sin^2x/cos^2x = tan^2x since tan x = sin x/cos x I'm not sure if this answers your question for an answer without sec

OpenStudy (anonymous):

thanks guys